# GazeMotion: Gaze-guided Human Motion Forecasting

### v1

@@ I Introduction
In parallel line of work, studies in the cognitive and behavioural sciences have revealed the strong correlation between human eye gaze and body movements during daily activities [8, 9]. For example, when navigating, human visual attention is strongly correlated with the movement direction [10] and it tends to precede corresponding hand and arm movements when reaching for an object [11]. Despite the close coordination between human eye gaze and body movements, this information has only recently started to being explored for human motion forecasting but the method proposed in [12] still required rich information about the full 3D environment and objects therein.
We present GazeMotion – the first learning-based method for human motion forecasting that combines information on past human poses with eye gaze without requiring such information. Our method consists of three main components: a convolutional neural network to predict future eye gaze from historical gaze, a pose-gaze graph that fuses pose and gaze data, as well as a novel residual graph convolutional network consisting of three spatial-temporal modules that forecasts future body poses from the pose-gaze graph. We evaluate our method for motion forecasting at different future time horizons of up to $1$ second (future 30 frames) on the MoGaze [13], ADT [14], as well as GIMO [12] benchmark datasets. We show that our method outperforms several state-of-the-art methods by a large margin: We achieve 7.4% improvement on MoGaze, 6.4% on ADT, and 6.1% on GIMO in terms of mean per joint position error (MPJPE). We also incorporate eye gaze into state-of-the-art methods and validate that our method still outperforms them by a large margin (5.6%), validating the superiority of our integration. We further report an online user study that shows that our method outperforms other methods in terms of precision and perceived realism of the predicted human motion. Considering that eye gaze information is not always available in real applications, we further use head direction as a proxy to gaze and show that our method still achieves significantly better performances than prior methods  [^1] Source code and trained models will be released upon acceptance..
@@
  • We report extensive experiments on three public datasets for motion forecasting at different future time horizons and demonstrate significant performance improvements over several state-of-the-art methods and conduct an online user study and validate that our method outperforms prior methods in both precision and realism.
  • We introduce to use head direction as a proxy to gaze in situations where eye gaze is not available to increase the usability of our method.
@@ II-A Coordination of Eye and Body Movements
The coordination of human eye and body movements has been extensively studied in multiple fields, including cognitive science and human-centred computing. Hu et al. [15, 16] have demonstrated that head movements are strongly correlated with eye movements in many daily activities, such as free viewing or searching for an object. Emery et al. [11] investigated the coordination of eye, hand, and head movements in a virtual environment. In this work, we demonstrate that findings on the strong link between human motion and eye gaze can be successfully transferred to the task of motion forecasting and lead to significant performance improvements.
@@ II-B Eye Gaze Prediction
Human eye gaze is significant for many important applications including human-robot interaction [1, 2] and human action anticipation [17]. In this background, eye gaze prediction has become a popular research topic in the areas of robotics and human-centred computing. Kim et al. focused on robot manipulation scenarios and proposed a Transformer-based method to predict eye gaze based on sequential visual input [18]. Hu et al. concentrated on a visual search task and used the scene content and task-related information to forecast future eye gaze [16]. Existing methods usually rely on additional information, e.g. image content and task-related variables, making it difficult to apply their methods to other situations. In contrast, our method only employs historical eye gaze to forecast future gaze. In addition, we are the first to combine gaze prediction with motion forecasting, i.e. we first forecast future eye gaze and then use the predicted future gaze to forecast future motion.
@@ III Method
We define gaze-guided human motion forecasting as the task of predicting future body motion, for a certain time horizon, jointly from past body poses and eye gaze information. Human pose $p$ is represented by the 3D positions of all human joints $p\in R^{3\times n}$, where $n$ is the number of joints, and employed a unit vector $g\in R^{3}$ to indicate eye gaze direction. Given a sequence of past human poses $P_{1:t}=\{p_{1},p_{2},...,p_{t}\}$ and human eye gaze $G_{1:t}=\{g_{1},g_{2},...,g_{t}\}$, the task is to forecast human poses in the future $P_{t+1:T}=\{p_{t+1},p_{t+2},...,p_{T}\}$. Our method consists of three main components: a gaze prediction network that forecasts future eye gaze from past gaze, a gaze-pose fusion procedure that fuses predicted future gaze and past poses into a gaze-pose graph, and a motion forecasting network that forecasts future motions from the fused gaze and motion data (see Figure 1 for an overview of our method).
@@ III-A Gaze Prediction
Guided by the intuition that information on future eye gaze is more useful for motion forecasting than past eye gaze, we first forecast future gaze $G_{t+1:2t}$ from past gaze $G_{1:t}$ and then use the predicted future gaze to forecast future body motion. In light of the good performance of 1D CNN for processing eye gaze data [19, 16], we employed four 1D CNN layers to forecast eye gaze. More specifically, we used three 1D CNN layers, each with $32$ channels and a kernel size of three, to extract features from the historical eye gaze data. Each CNN layer was followed by a layer normalisation (LN) and a Tanh activation function. After the three CNN layers, we used a 1D CNN layer with three channels and a kernel size of three, and a Tanh activation function to predict future eye gaze from the gaze features. The predicted gaze were finally normalised to unit vectors: $\hat{G}_{t+1:2t}=\{\hat{g}_{t+1},\hat{g}_{t+2},...,\hat{g}_{2t}\}\in R^{3\times t}$.
@@ III-B Gaze-Pose Fusion
We first padded the past poses and predicted gaze up to $T$ by repeating the last pose and gaze for $T-t$ times following prior works [6, 20]. Each pose was represented using the 3D coordinates of all the human joints: $P_{1:T}=\{J^{1}_{1:T},J^{2}_{1:T},...,J^{n}_{1:T}\}\in R^{3\times n\times T}$. We further cloned the predicted eye gaze data $\hat{G}$ for $n-1$ times to make the gaze data have the same size as the motion data. We further concatenated the motion data and gaze data along the spatial dimension and obtained $X\in R^{3\times 2n\times T}$. We modelled the fused motion and gaze data $X$ as fully-connected spatial and temporal graphs, jointly called the gaze-pose graph, to learn the relationship between human eye gaze and human motion. The spatial graph represents each joint data $J^{k}$ and each gaze data $\hat{G}$ as a separate node and thus contains $2n$ nodes: $n$ joint nodes and $n$ gaze nodes. The temporal graph consists of $T$ nodes, corresponding to the fused data at different times: $X_{1},X_{2},...,X_{T}$. The spatial and temporal graphs are both fully-connected with their adjacency matrices measuring the weights between each pair of nodes.
@@ III-C Motion Forecasting
We then proposed three novel GCN modules, i.e. a start module, a residual module, and an end module, to process the transformed data $X_{d}\in R^{3\times 2n\times T}$.
@@
Residual Module: The residual module consists of $m$ GCN blocks with each block containing a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{2n\times 2n}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to prevent from overfitting. $m$ was set to $16$ and a residual connection was added for each GCN block. We cut the output of the residual module in half in the temporal dimension ($R^{16\times 2n\times 2T}\rightarrow R^{16\times 2n\times T}$) and input it to the end module.
End Module: The end module consists of a temporal GCN that learns the temporal adjacency matrix, a weight matrix $W^{end}\in R^{16\times 3}$ that maps the latent features to $3$ dimensions, and a spatial GCN that learns the spatial adjacency matrix. We added a global residual connection to improve the network flow. The output of the end module $Y_{d}\in R^{3\times 2n\times T}$ was converted back to the original representation space using an inverse discrete cosine transform (IDCT) matrix $M_{idct}\in R^{T\times T}$:
@@ IV-A Datasets
The ADT dataset contains human pose and gaze data performing various indoor activities including room decoration, meal preparation, and work. Each human pose consists of the 3D coordinates of $21$ human joints recorded at $30$ Hz. For experiments on ADT, we randomly selected $25$ sequences for training and $10$ sequences for testing.
@@ IV-B Evaluation Settings
For experiments on the MoGaze, ADT, and GIMO datasets (30 Hz), we used $10$ frames of data as input to forecast human poses in the future $30$ frames (i.e. up to one second into the future), following the common evaluation settings  [6, 5].
@@ IV-C Motion Forecasting Results
IV-D Adding Eye Gaze to Other Methods
Existing motion forecasting methods ignored the effectiveness of eye gaze for predicting human motions. To make a more comprehensive comparison, we also added eye gaze to existing methods using our gaze-pose fusion strategy. Table II shows the performances of each method for motion forecasting on the GIMO dataset. We find that our method consistently outperforms the state of the art in terms of performances at different time intervals as well as average performance, achieving an average improvement of 5.6% ($83.8$ vs. $88.8$). Paired Wilcoxon signed-rank test validates that the differences between our method and the state of the art are statistically significant ($p<0.01$).
TABLE II: MPJPE errors (unit: millimeters) of adding eye gaze to different methods for motion forecasting on GIMO.
Method 200 ms 400 ms 600 ms 800 ms 1000 ms Average
Res-RNN [4] + gaze 81.5 122.5 164.1 205.7 246.6 148.4
siMLPe [7] + gaze 43.0 77.0 112.6 147.2 182.6 98.4
HisRep [5] + gaze 45.3 82.2 118.1 154.2 190.5 102.5
PGBIG [6] + gaze 36.9 69.3 101.9 135.5 169.5 88.8
Ours 32.6 64.1 97.0 130.0 162.4 83.8
IV-E Head Direction as a Proxy to Eye Gaze
@@ IV-F Ablation Study
We tested different ways of using eye gaze information, i.e. 1) don’t use eye gaze (w/o gaze), 2) use past head direction in the gaze-pose fusion procedure (past head), 3) use predicted future head direction (future head), 4) use past eye gaze (past gaze), and 5) use predicted future eye gaze (Ours). Table III shows the motion forecasting results of different ways of using eye gaze on the MoGaze dataset. As can be seen from the table, using eye gaze or head direction as a proxy to eye gaze achieves significantly better performances than not using eye gaze (paired Wilcoxon signed-rank test, $p<0.01$), validating that eye gaze information can help improve the performance of motion forecasting. We also find that using predicted future eye gaze or head direction achieves better performances than directly using past eye gaze ($75.9$ vs. $77.1$) or head direction ($77.5$ vs. $77.8$), revealing the significant potential of applying gaze prediction methods to improve the performances of motion forecasting.
TABLE III: Ablation study on the MoGaze dataset.
@@
From Table I we can see that even without using eye gaze, our method still significantly outperforms the state of the art ($79.0$ vs. $82.0$ on MoGaze, $39.1$ vs. $40.6$ on ADT, $86.8$ vs. $89.2$ on GIMO, paired Wilcoxon signed-rank test, $p<0.01$), validating the effectiveness of our GCN architecture. We also performed an ablation study to evaluate the effectiveness of each component in our method. Specifically, we respectively removed spatial GCN, temporal GCN, global residual, and velocity loss, and retrained the ablated methods. We can see from Table III that each component helps improve our method’s motion forecasting performance.
IV-G User Study
@@
IV-G1 Stimuli
We randomly selected $24$ motion forecasting samples from the MoGaze, ADT, and GIMO datasets ($8$ samples from one dataset) and used them as our stimuli. Each sample consisted of $30$ frames of predictions (corresponding to future $1$ second) and was visualised as a short video.
IV-G2 Participants
@@ IV-G3 Procedure
IV-G4 Statistical Analysis
The means and standard deviations (SDs) of different methods’ rankings are shown in Table IV. We can see that our method outperforms the state of the art in terms of both precision ($1.6$ vs. $3.2$) and realism ($1.9$ vs. $3.1$) and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$). The above results demonstrate that our method achieves significantly better performances over prior methods in qualitative evaluation.
TABLE IV: Statistical results of different methods’ rankings in our user study. Best results are in bold while second best are underlined.

### v2

@@ I Introduction
In a parallel line of work, studies in the cognitive and behavioural sciences have revealed the strong correlation between human eye gaze and body movements during daily activities [8, 9]. For example, when navigating, human visual attention is strongly correlated with the movement direction [10] and it tends to precede corresponding hand and arm movements when reaching for an object [11]. Despite the close coordination between human eye gaze and body movements, this information has only recently started to be explored for human motion forecasting. A method proposed in [12] requires rich information about the full 3D environment and objects therein.
We present GazeMotion – the first learning-based method for human motion forecasting that combines information on past human poses with eye gaze without requiring such information. Our method consists of three main components: a convolutional neural network to predict future eye gaze from historical gaze, a pose-gaze graph that fuses pose and gaze data, and a novel residual graph convolutional network consisting of three spatial-temporal modules that forecasts future body poses from the pose-gaze graph. We evaluate our method for motion forecasting at different future time horizons of up to $1$ second (future 30 frames) on the MoGaze [13], ADT [14], and GIMO [12] benchmark datasets. We show that our method outperforms several state-of-the-art methods by a large margin: We achieve 7.4% improvement on MoGaze, 6.4% on ADT, and 6.1% on GIMO in terms of mean per joint position error (MPJPE). We further report an online user study that shows that our method outperforms other methods in terms of precision and perceived realism of the predicted human motion. Considering that eye gaze information is not always available in real applications, we further use head direction as a proxy to gaze and show that our method still achieves significantly better performances than prior methods. The full source code and trained models are available at zhiminghu.net/hu24_gazemotion.
@@
  • We report extensive experiments on three public datasets for motion forecasting at different future time horizons and demonstrate significant performance improvements over several state-of-the-art methods.
  • We conduct an online user study and validate that our method outperforms prior methods in both precision and realism.
@@ II-A Coordination of Eye and Body Movements
The coordination of human eye and body movements has been extensively studied in multiple fields, including cognitive science and human-centred computing. Hu et al. [15, 16, 17] have demonstrated that head movements are strongly correlated with eye movements in many daily activities, such as free viewing or searching for an object. Emery et al. [11] investigated the coordination of eye, hand, and head movements in a virtual environment. In this work, we demonstrate that findings on the strong link between human motion and eye gaze can be successfully transferred to the task of motion forecasting and lead to significant performance improvements.
@@ II-B Eye Gaze Prediction
Human eye gaze is significant for many important applications including human-robot interaction [1, 2] and human action anticipation [18, 19]. In this background, eye gaze prediction has become a popular research topic in the areas of robotics and human-centred computing. Kim et al. focused on robot manipulation scenarios and proposed a Transformer-based method to predict eye gaze based on sequential visual input [20]. Hu et al. concentrated on a visual search task and used the scene content and task-related information to forecast future eye gaze [16]. Existing methods usually rely on additional information, e.g. image content and task-related variables, making it difficult to apply their methods to other situations. In contrast, our method only employs historical eye gaze to forecast future gaze. In addition, we are the first to combine gaze prediction with motion forecasting, i.e., we first forecast future eye gaze and then use the predicted future gaze to forecast future motion.
@@ III Method
We define gaze-guided human motion forecasting as the task of predicting future body motion, for a certain time horizon, jointly from past body poses and eye gaze information. Human pose $p$ is represented by the 3D positions of all human joints $p\in R^{3\times n}$, where $n$ is the number of joints. Eye gaze direction is indicated using a unit vector $g\in R^{3}$. Given a sequence of past human poses $P_{1:t}=\{p_{1},p_{2},...,p_{t}\}$ and human eye gaze $G_{1:t}=\{g_{1},g_{2},...,g_{t}\}$, the task is to forecast human poses in the future $P_{t+1:T}=\{p_{t+1},p_{t+2},...,p_{T}\}$. Our method consists of three main components: a gaze prediction network that forecasts future eye gaze from past gaze, a gaze-pose fusion procedure that fuses predicted future gaze and past poses into a gaze-pose graph, and a motion forecasting network that forecasts future motions from the fused gaze and motion data (see Figure 1 for an overview of our method).
@@ III-A Gaze Prediction
Guided by the intuition that information on future eye gaze is more useful for motion forecasting than past eye gaze, we first forecast future gaze $G_{t+1:2t}$ from past gaze $G_{1:t}$ and then use the predicted future gaze to forecast future body motion. In light of the good performance of 1D CNN for processing eye gaze data [21, 16], we employed four 1D CNN layers to forecast eye gaze. More specifically, we used three 1D CNN layers, each with $32$ channels and a kernel size of three, to extract features from the historical eye gaze data. Each CNN layer was followed by a layer normalisation (LN) and a Tanh activation function. After the three CNN layers, we used a 1D CNN layer with three channels and a kernel size of three, and a Tanh activation function to predict future eye gaze from the gaze features. The predicted gaze directions were finally normalised to unit vectors: $\hat{G}_{t+1:2t}=\{\hat{g}_{t+1},\hat{g}_{t+2},...,\hat{g}_{2t}\}\in R^{3\times t}$.
@@ III-B Gaze-Pose Fusion
We first padded the past poses and predicted gaze up to $T$ by repeating the last pose and gaze for $T-t$ times following prior works [6, 22]. Each pose was represented using the 3D coordinates of all the human joints: $P_{1:T}=\{J^{1}_{1:T},J^{2}_{1:T},...,J^{n}_{1:T}\}\in R^{3\times n\times T}$. We further cloned the predicted eye gaze data $\hat{G}$ for $n-1$ times to make the gaze data have the same size as the motion data. We then concatenated the motion data and gaze data along the spatial dimension and obtained $X\in R^{3\times 2n\times T}$. We modelled the fused motion and gaze data $X$ as fully-connected spatial and temporal graphs, jointly called the gaze-pose graph, to learn the relationship between human eye gaze and human motion. The spatial graph represents each joint $J^{k}$ and each gaze $\hat{G}$ as a separate node and thus contains $2n$ nodes: $n$ joint nodes and $n$ gaze nodes. The temporal graph consists of $T$ nodes, corresponding to the fused data at different times: $X_{1},X_{2},...,X_{T}$. The spatial and temporal graphs are both fully-connected with their adjacency matrices measuring the weights between each pair of nodes.
@@ III-C Motion Forecasting
We then proposed three novel GCN modules, i.e., a start module that mapped $X_{d}\in R^{3\times 2n\times T}$ into feature space, a residual module that extracted the features, and an end module that mapped the features to the original data space.
@@
Residual Module: The residual module consisted of $m$ GCN blocks with each block containing a temporal GCN that learned the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracted the latent features, a spatial GCN that learned the spatial adjacency matrix $A^{S}_{i}\in R^{2n\times 2n}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to prevent overfitting. $m$ was set to $16$ and a residual connection was added for each GCN block. We cut the output of the residual module in half in the temporal dimension ($R^{16\times 2n\times 2T}\rightarrow R^{16\times 2n\times T}$) and input it to the end module.
End Module: The end module consisted of a temporal GCN that learned the temporal adjacency matrix, a weight matrix $W^{end}\in R^{16\times 3}$ that mapped the latent features to $3$ dimensions, and a spatial GCN that learned the spatial adjacency matrix. We added a global residual connection to improve the network flow. The output of the end module $Y_{d}\in R^{3\times 2n\times T}$ was converted back to the original representation space using an inverse discrete cosine transform (IDCT) matrix $M_{idct}\in R^{T\times T}$:
@@ IV-A Datasets
The ADT dataset contains $35$ sequences of human pose and gaze data performing various indoor activities including room decoration, meal preparation, and work. Each human pose consists of the 3D coordinates of $21$ human joints recorded at $30$ Hz. For experiments on ADT, we randomly selected $25$ sequences for training and $10$ sequences for testing.
@@ IV-B Evaluation Settings
For experiments on the MoGaze, ADT, and GIMO datasets (30 Hz), we used $10$ frames of data as input to forecast human poses in the future $30$ frames (i.e., up to one second into the future), following the common evaluation settings [6, 5].
@@ IV-C Motion Forecasting Results
IV-D Head Direction as a Proxy to Eye Gaze
@@ IV-E Ablation Study
We tested different ways of using eye gaze information, i.e., 1) don’t use eye gaze (w/o gaze), 2) use past head direction in the gaze-pose fusion procedure (past head), 3) use predicted future head direction (future head), 4) use past eye gaze (past gaze), and 5) use predicted future eye gaze (Ours). Table II shows the motion forecasting results of different ways of using eye gaze on the MoGaze dataset. As can be seen from the table, using eye gaze or head direction as a proxy to eye gaze achieves significantly better performances than not using eye gaze (paired Wilcoxon signed-rank test, $p<0.01$), validating that eye gaze information can help improve the performance of motion forecasting. We also find that using predicted future eye gaze or head direction achieves better performances than directly using past eye gaze ($75.9$ vs. $77.1$) or head direction ($77.5$ vs. $77.8$), revealing the significant potential of applying gaze prediction methods to improve the performances of motion forecasting.
TABLE II: Ablation study on the MoGaze dataset.
@@
From Table I we can see that even without using eye gaze, our method still significantly outperforms the state of the art ($79.0$ vs. $82.0$ on MoGaze, $39.1$ vs. $40.6$ on ADT, $86.8$ vs. $89.2$ on GIMO, paired Wilcoxon signed-rank test, $p<0.01$), validating the effectiveness of our GCN architecture. Furthermore, we respectively removed spatial GCN, temporal GCN, copy in start module, global residual, and velocity loss, and retrained the ablated methods. We can see from Table II that each component helps improve our method’s motion forecasting performance. We also find that temporal GCN is most important to the success of our method, revealing the significance of temporal features for motion forecasting.
Training Parameters
We used different dropout rate to train our model and the performances of dropout rate $0.1$, $0.2$, $0.3$ (Ours), and $0.4$ on MoGaze are $77.7$, $76.9$, $75.9$, and $76.5$, respectively. We also added a weighting constant for $\ell_{v}$ in Equation 7 and the performances of weighting constant $0.25$, $0.5$, $1.0$ (Ours), and $1.5$ on MoGaze are $76.0$, $76.5$, $75.9$, and $76.3$, respectively. These results validate that our training parameters are optimal in practice.
IV-F User Study
@@
IV-F1 Stimuli
We randomly selected $24$ motion forecasting samples from the MoGaze, ADT, and GIMO datasets ($8$ samples from each dataset) and used them as our stimuli. Each sample consisted of $30$ frames of predictions (corresponding to future $1$ second) and was visualised as a short video.
IV-F2 Participants
@@ IV-F3 Procedure
IV-F4 Statistical Analysis
The means and standard deviations (SDs) of different methods’ rankings are shown in Table III. We can see that our method outperforms the state of the art in terms of both precision ($1.6$ vs. $3.2$) and realism ($1.9$ vs. $3.1$) and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$). The above results demonstrate that our method achieves significantly better performances over prior methods in qualitative evaluation.
TABLE III: Statistical results of different methods’ rankings in our user study.
@@ IV-F4 Statistical Analysis
IV-G Discussion
From the results in Table I, we find that the performances of all the methods deteriorate significantly with the increase of prediction time. This is a well-known problem in motion forecasting [6, 7] that all current methods suffer from since these methods only use historical motion information. Integrating more context information such as user’s goal or task into motion forecasting has the potential to alleviate this problem. In addition, we only explored the effectiveness of eye gaze and head direction on motion forecasting but ignored other important body signals such as hand gestures or gait. Integrating such body signals into our pipeline to further improve the performance is an interesting avenue of future work. Furthermore, we are also looking forward to incorporating stochasticity into the human motion forecasting model to further improve the performance.